# E-scooter effects on public transport demand: a case study in Santiago, Chile

### v1

@@ Appendix C Results of Negative Binomial Models for Evaluating Generation of Public Transport Trips
Log(working population) -0.182 (-13.377) *** 0.202 (7.693) ***
Relative educational level 0.255 (23.487) *** 0.411 (21.753) ***
Model performance
Nº observations 20,079 7,996
Log-likelihood -1.3852e+05 -61517
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 12: Results of negative binomial model for evaluating generation of metro trips
Parameters Intermediate Central
Region Region
Intercept -93.041 (-13.088) *** 183.3205 (17.157) ***
Morning peak 178.7567 (11.575) *** -47.9854 (-2.283) **
Lunch -50.9498 (-3.161) ** 114.8304 (5.053) ***
Afternoon peak 165.2241 (10.947) *** 350.3334 (16.208) ***
Afternoon valley 110.6648 (7.406) *** 286.2683 (13.460) ***
Night -1447.3674 (-57.926) *** -1690.5511 (-55.971) ***
Night valley -289.2255 (-17.965) *** -214.8078 (-9.916) ***
Friday 33.9532 (2.748) ** 27.0006 (1.668) *
Saturday -244.4253 (-18.360) *** -326.0656 (-18.950) ***
Sunday -459.9325 (-30.774) *** -652.9032 (-33.563) ***
Year 2019 -46.5152 (-5.174) *** 12.1303 (1.087)
Scooter trip destination 365.3049 (16.333) *** -136.6634 (-2.870) **
Metro -93.0410 (-13.088) *** 119.1737 (5.758) ***
Log(working population) 183.3205 (17.157) *** 183.3205 (17.157) ***
Relative educational level 130.5021 (22.234) *** 344.3689 (36.405) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 13: Marginal effects of negative binomial models for evaluating generation of metro trips
@@ Appendix D Results of Negative Binomial Models for Evaluating Attraction of Public Transport Trips
Morning peak 0.284 (21.874) *** 0.148 (11.308) *** 0.159 (4.707) ***
Lunch 0.221 (16.331) *** 0.015 (1.097) -0.166 (-4.604) ***
Afternoon peak 0.996 (79.981) *** 0.409 (32.227) *** -0.081 (-2.454) *
Afternoon valley 0.631 (51.610) *** 0.267 (21.396) *** -0.043 (-1.313)
Night -1.420 (-100.531) *** -2.207 (-165.063) *** -3.025 (-96.222) ***
Night valley 0.328 (24.474) *** -0.378 (-27.806) *** -1.019 (-28.876) ***
@@
Friday 0.024 (2.346) ** 0.038 (3.735) *** 0.041 (1.611)
Saturday -0.482 (-45.596) *** -0.488 (-46.464) *** -0.533 (-20.183) ***
Sunday -0.919 (-83.369) *** -0.930 (-85.292) *** -1.071 (-39.088) ***
Treatment period
Year 2019 0.010 (1.408) -0.046 (-6.388) *** -0.085 (-2.714) ***
Treatment zones
Scooter trips origins 0.188 (1.425) 0.654 (26.747) *** -0.762 (-9.003) ***
@@
Log(working population) 0.075 (17.477) *** 0.268 (48.505) *** 0.317 (13.165) ***
Relative educational level 0.165 (34.366) *** 0.278 (63.178) *** 0.602 (49.025) ***
@@
Table 14: Results of negative binomial models for evaluating attraction of metro and bus trips
@@
Morning peak 19.716 (1.745) *** 81.130 (17.283) *** 20.006 (0.920) ***
Lunch 2.019 (1.840) -84.628 (18.433) *** 15.592 (0.959) ***
Afternoon peak 54.549 (1.736) *** -41.443 (16.904) ** 70.270 (0.991) ***
Afternoon valley 35.610 (1.684) *** -21.880 (16.662) 44.500 (0.906) ***
Night -294.399 (2.366) *** -1544.635 (26.621) *** -100.253 (1.115) ***
Night valley -50.349 (1.833) *** -520.130 (19.738) *** 23.117 (0.955) ***
Year 2019 -6.067 (0.950) *** 24.765 (9.131) ** 0.699 (0.496)
Friday 5.067 (1.356) *** 21.203 (13.166) 1.678 (0.715) **
Saturday -65.059 (1.453) *** -272.559 (14.120) *** -34.027 (0.772) ***
Sunday -124.000 (1.623) *** -547.100 (16.052) *** -64.856 (0.861) ***
Scooter trips origins 87.171 (3.346) *** -389.172 (43.670) *** 13.275 (9.314)
Metro 8.695 (0.971) *** 61.877 (16.794) *** 5.319 (0.305) ***
Log(working population) 35.683 (0.759) *** 161.874 (9.187) *** 11.659 (0.349) ***
Relative educational level 37.051 (0.638) *** 307.614 (7.839) *** 11.659 (0.349) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 15: Marginal effects of negative binomial models for evaluating attraction of bus and metro trips
@@
Morning peak 0.291 (25.449) *** 0.141 (13.393) *** 0.155 (5.860) ***
Lunch 0.192 (16.077) *** -0.009 (-0.777) -0.170 (-5.971) ***
Afternoon peak 0.866 (78.818) *** 0.292 (28.503) *** -0.211 (-8.137) ***
Afternoon valley 0.561 (51.997) *** 0.215 (21.315) *** -0.046 (-1.803)
Night -1.271 (-101.055) *** -1.818 (-166.492) *** -2.199 (-88.225) ***
Night valley 0.234 (19.854) *** -0.438 (-39.816) *** -1.180 (-42.493) ***
@@
Friday 0.008 (0.918) 0.027 (3.261) *** 0.034 (1.620)
Saturday -0.477 (-51.174) *** -0.453 (-53.279) *** -0.434 (-20.779) ***
Sunday -0.875 (-89.852) *** -0.859 (-97.042) *** -0.887 (-40.966) ***
Treatment period
Year 2019 -0.036 (-5.853) *** -0.065 (-11.305) *** -0.187 (-13.207) ***
Treatment zones
Scooter trips origins 0.789 (5.829) *** 0.225 (11.543) *** 0.391 (5.897) ***
@@
Log(working population) 0.246 (55.428) *** 0.354 (75.494) *** 0.157 (12.710) ***
Relative educational level 0.060 (15.574) *** 0.184 (63.178) *** 0.336 (33.561) ***
@@
Table 16: Results of negative binomial models for evaluating attraction of bus trips
@@
Morning peak 14.253 (0.564) *** 8.779 (0.657) *** 20.363 (3.485) ***
Lunch 9.403 (0.587) *** -0.541 (0.695) -22.236 (3.736) ***
Afternoon peak 42.421 (0.589) *** 18.159 (0.645) *** -27.713 (3.426) ***
Afternoon valley 27.473 (0.549) *** 13.374 (0.632) *** -6.057 (3.360)
Night -62.248 (0.674) *** -113.066 (0.832) *** -288.239 (4.559) ***
Night valley 11.502 (0.582) *** -27.219 (0.696) *** -154.566 (4.074) ***
Year 2019 -1.772 (0.303) *** -4.051 (0.359) *** -24.431 (1.878) ***
Friday 0.401 (0.437) 1.678 (0.512) *** 4.383 (2.660)
Saturday -23.383 (0.472) *** -28.196 (0.546) *** -56.828 (2.811) ***
Sunday -42.875 (0.523) *** -53.444 (0.603) *** -116.202 (3.112) ***
Scooter trips origins 38.622 (5.662) *** 13.996 (1.217) *** 51.162 (8.711) ***
Metro 7.073 (0.332) *** -2.502 (0.370) *** 54.786 (3.524) ***
Log(working population) 12.062 (1.492) *** 22.048 (0.311) *** 12.062 (1.644) ***
Relative educational level 2.930 (0.189) *** 11.467 (0.224) *** 44.086 (1.422) ***
@@
Table 17: Marginal effects of negative binomial models for evaluating attraction of bus trips
@@
Morning peak 0.262 (9.040) *** 0.224 (4.848) ***
Lunch 0.033 (1.072) -0.174 (-3.577) ***
Afternoon peak 0.543 (19.337) *** -0.038 (-0.849)
Afternoon valley 0.350 (12.652) *** -0.025 (-0.565)
Night -2.371 (-71.668) *** -3.070 (-60.407) ***
Night valley -0.319 (-10.749) *** -0.985 (-20.666) ***
Treatment period
Year 2019 -0.061 (-4.076) *** 0.045 (1.786)
@@
Friday 0.066 (2.468) ** 0.044 (1.198)
Saturday -0.553 (-23.068) *** -0.637 (-16.603) ***
Sunday -0.978 (-38.578) *** -1.173 (-28.528) ***
Treatment zones
Scooter trips origins 0.627 (14.874) *** 0.395 (1.408)
@@
Log(working population) -0.175 (-13.360) *** 0.089 (3.253) ***
Relative educational level 0.329 (32.655) *** 0.482 (25.921) ***
Model performance
Nº observations 24,579 9,283
Log-likelihood -1.6604e+05 -70503
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 18: Results of negative binomial models for evaluating attraction of metro trips
Parameters Intermediate Central
Region Region
Morning peak 119.811 (13.318) *** 259.532 (53.871) ***
Lunch 14.836 (13.838) -201.826 (56.595) ***
Afternoon peak 247.817 (13.417) *** -44.194 (52.047)
Afternoon valley 159.727 (12.879) *** -29.187 (51.691)
Night -1082.687 (19.720) *** -3560.477 (89.221) ***
Night valley -145.858 (13.658) *** -1142.470 (60.046) ***
Year 2019 -30.495 (7.489) *** 52.561 (29.450)
Friday 25.798 (10.462) ** 50.876 (42.487)
Saturday -252.478 (11.465) *** -738.275 (47.236) ***
Sunday -446.582 (12.996) *** -1361.016 (55.188) ***
Scooter trips origins 286.414 (19.966) *** 458.374 (325.778)
Log(working population) -79.894 (6.110) *** 102.957 (31.733) ***
Relative educational level 150.112 (5.106) *** 559.393 (24.861) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 19: Marginal effects of negative binomial models for evaluating attraction of metro trips

### v2

@@ Appendix C Results of Negative Binomial Models for Evaluating Generation of Public Transport Trips
Log(working population) -243.0255 (-6.675) *** 950.9396 (7.768) ***
Relative educational level 395.3644 (10.053) *** 1836.4571 (18.737) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 15: Marginal effects of negative binomial models for evaluating generation of metro trips by region.
@@ Appendix D Results of Negative Binomial Models for Evaluating Attraction of Public Transport Trips
Morning peak 0.2828 (21.819) *** 0.1537 (11.730) *** 0.1627 (4.822) ***
Lunch 0.2202 (16.270) *** 0.0098 (0.711) -0.1602 (-4.620) ***
Afternoon peak 0.9948 (79.825) *** 0.3999 (31.433) *** -0.0851 (-2.575) **
Afternoon valley 0.6297 (51.501) *** 0.2616 (20.915) *** -0.0434 (-1.331)
Night -1.4215 (-100.573) *** -2.2359 (-167.240) *** -3.0218 (-96.218) ***
Night valley 0.3266 (24.384) *** -0.3893 (-28.602) *** -1.0225 (-28.996) ***
@@
Friday 0.0237 (2.339) ** 0.0383 (3.757) *** 0.0430 (1.669) *
Saturday -0.4823 (-45.604) *** -0.4905 (-46.614) *** -0.5355 (-20.258) ***
Sunday -0.9193 (-83.384) *** -0.9338 (-85.465) *** -1.0701 (-39.063) ***
Difference-in-differences variables
Post 0.0098 (1.384) -0.0099 (-1.306) 0.0378 (2.070) **
Treatment_zones -0.0785 (-1.521) -0.1131 (-6.183) *** -0.4432 (-7.702) ***
Treatment_zones x post 0.0436 (0.613) 0.0418 (2.048) ** -0.0306 (-0.390)
@@
Log(working population) 0.0752 (17.429) *** 0.2806 (50.788) *** 0.3079 (18.051) ***
Relative educational level 0.1681 (33.372) *** 0.3412 (60.611) *** 0.6298 (49.526) ***
@@
Table 16: Results of negative binomial models for evaluating attraction of bus and metro trips.
@@
Time periods
Morning peak 19.9543 (21.684) *** 20.4836 (11.718) *** 82.9149 (4.809) ***
Lunch 15.5340 (16.190) *** 1.3097 (0.711) -84.7186 (-4.607) ***
Afternoon peak 70.1871 (70.766) *** 53.2886 (30.724) *** -43.3861 (-2.572) **
Afternoon valley 44.4302 (49.047) *** 34.8545 (20.698) *** -22.1223 (-1.331)
Night -100.2964 (-89.925) *** -297.9469 (-125.879) *** -1540.2396 (-58.082) ***
Night valley 23.0469 (24.123) *** -51.8750 (-28.229) *** -521.1965 (-26.443) ***
Type of day
Friday 1.6727 (2.339) ** 5.0994 (3.756) *** 21.9143 (1.668) *
Saturday -34.0297 (-44.065) *** -65.3573 (-44.930) *** -272.9721 (-19.368) ***
Sunday -64.8611 (-75.347) *** -124.4369 (-76.670) *** -545.4409 (-34.068) ***
Difference-in-differences variables
Post 0.6898 (1.384) -1.3164 (-1.306) 19.2547 (2.069) **
Treatment zones -5.5352 (-1.521) -15.0703 (-6.178) *** -225.8979 (-7.642) ***
Treatment zones × post 3.0729 (0.613) 5.5681 (2.048) ** -15.5797 (-0.390)
Infrastructure
Metro 12.6630 (22.830) *** 13.7034 (13.570) *** 67.0650 (3.996) ***
Sociodemographic information
Log(working population) 5.3062 (17.402) *** 37.3967 (49.008) *** 156.9238 (17.158) ***
Relative educational level 11.8625 (32.499) *** 45.4711 (56.300) *** 320.9919 (39.655) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 17: Marginal effects of negative binomial models for evaluating attraction of metro and bus trips by region.
@@
Morning peak 0.2968 (25.964) *** 0.1404 (13.320) *** 0.1534 (5.780) ***
Lunch 0.1959 (16.421) *** -0.0103 (-0.922) -0.1692 (-5.949) ***
Afternoon peak 0.8740 (79.589) *** 0.2905 (28.351) *** -0.2110 (-8.112) ***
Afternoon valley 0.5680 (52.686) *** 0.2135 (21.153) *** -0.0448 (-1.746) *
Night -1.2645 (-100.594) *** -1.8244 (-167.301) *** -2.2068 (-88.479) ***
Night valley 0.2430 (20.555) *** -0.4401 (-40.027) *** -1.1807 (-42.503) ***
@@
Friday 0.0081 (0.908) 0.0271 (3.291) *** 0.0333 (1.640)
Saturday -0.4765 (-51.119) *** -0.4536 (-53.056) *** -0.4341 (-20.785) ***
Sunday -0.8750 (-89.882) *** -0.8604 (-97.142) *** -0.8876 (-40.977) ***
Difference-in-differences variables
Post -0.0358 (-5.777) *** -0.0564 (-9.221) *** -0.1808 (-12.515) ***
Treatment_zones 0.5419 (12.060) *** -0.0773 (-5.462) *** -0.1088 (-2.373) **
Treatment_zones x post 0.0867 (1.396) 0.0211 (1.276) 0.1060 (1.721) *
@@
Log(working population) 0.2484 (66.134) *** 0.3563 (75.823) *** 0.1387 (11.094) ***
Relative educational level 0.0358 (8.836) *** 0.2135 (48.373) *** 0.3555 (33.653) ***
@@
Table 18: Results of negative binomial models for evaluating attraction of bus trips.
@@
Time periods
Morning peak 14.5498 (25.757) *** 8.7350 (13.299) *** 20.0847 (5.763) ***
Lunch 9.6039 (16.359) *** -0.6415 (-0.922) -22.1579 (-5.931) ***
Afternoon peak 42.8414 (72.624) *** 18.0721 (28.025) *** -27.6286 (-8.064) ***
Afternoon valley 27.8394 (50.678) *** 13.2786 (21.012) *** -5.8653 (-1.745) *
Night -61.9803 (-92.099) *** -113.4862 (-136.316) *** -289.0129 (-63.333) ***
Night valley 11.9088 (20.441) *** -27.3759 (-39.330) *** -154.6264 (-37.946) ***
Type of day
Friday 0.3968 (0.908) 1.6837 (3.290) *** 4.3616 (1.640)
Saturday -23.3565 (-49.445) *** -28.2178 (-51.683) *** -56.8494 (-20.223) ***
Sunday -42.8876 (-82.027) *** -53.5216 (-88.691) *** -116.2411 (-37.348) ***
Difference-in-differences variables
Post -1.7549 (-5.767) *** -3.5092 (-9.209) *** -23.6747 (-12.353) ***
Treatment zones 26.5597 (12.023) *** -4.8109 (-5.460) *** -14.2478 (-2.372) **
Treatment zones x post 4.2511 (1.396) 1.3094 (1.276) 13.8759 (1.720) *
Infrastructure
Metro 7.1517 (21.563) *** -1.5514 (-4.101) *** 54.7859 (15.533) ***
Sociodemographic information
Log(working population) 12.1759 (63.498) *** 22.1658 (71.042) *** 18.1703 (10.864) ***
Relative educational level 1.7545 (8.826) *** 13.2804 (46.836) *** 46.5557 (31.138) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 19: Marginal effects of negative binomial models for evaluating attraction of bus trips by region.
Parameters Intermediate Region Central Region
Intercept 7.4561 (68.157) *** 6.0602 (25.746) ***
Time periods
Morning peak 0.2481 (8.567) *** 0.2224 (4.820) ***
Lunch 0.0360 (1.193) -0.1746 (-3.590) ***
Afternoon peak 0.5623 (20.076) *** -0.0421 (-0.939)
Afternoon valley 0.3622 (13.131) *** -0.0265 (-0.595)
Night -2.4647 (-75.216) *** -3.0731 (-60.502) ***
Night valley -0.3028 (-10.215) *** -0.9904 (-20.778) ***
Type of day
Friday 0.0587 (2.569) ** 0.0435 (1.188)
Saturday -0.5464 (-22.864) *** -0.6368 (-16.616) ***
Sunday -0.9734 (-38.500) *** -1.1742 (-28.558) ***
Difference-in-differences variables
Post 0.0225 (1.317) 0.0473 (1.821) *
Treatment_zones 0.7056 (17.541) *** -0.2021 (-2.332) **
Treatment_zones x post -0.0769 (-1.732) * 0.0079 (0.066)
Sociodemographic information
Log(working population) -0.1866 (-14.286) *** 0.0851 (3.116) ***
Relative educational level 0.2042 (15.812) *** 0.4994 (26.150) ***
Model performance
Nº observations 24,579 9,283
Log-likelihood -1.6595e+05 -7.0499e+04
@@
Table 20: Results of negative binomial models for evaluating attraction of metro trips.
@@
Morning peak 114.5955 (8.534) *** 257.9006 (4.791) ***
Lunch 16.6482 (1.192) -202.4349 (-3.579) ***
Afternoon peak 259.7544 (19.091) *** -48.8484 (-0.939)
Afternoon valley 167.3319 (12.848) *** -30.7297 (-0.595)
Night -1138.6288 (-55.833) *** -3563.4397 (-39.943) ***
Night valley -139.8859 (-10.162) *** -1148.4773 (-19.108) ***
@@
Friday 27.1079 (2.567) ** 50.4060 (1.187)
Saturday -252.4360 (-21.858) *** -738.4166 (-15.640) ***
Sunday -449.6835 (-34.309) *** -1361.5697 (-24.684) ***
Difference-in-differences variables
Post 10.3876 (1.317) 54.7903 (1.820) *
Treatment zones 325.9625 (16.743) *** -234.3963 (-2.328) **
Treatment zones x post -35.5166 (-1.732) * 9.1316 (0.066)
@@
Log(working population) -86.2191 (-13.930) *** 98.6420 (3.108) ***
Relative educational level 94.3309 (15.452) *** 579.0490 (22.690) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parentheses.
Table 21: Marginal effects of negative binomial models for evaluating attraction of metro trips by region.